OpenStudy (anonymous):

hich has the greatest value? zero to the fourth power 2 to the fourth power seven to the second power ten to the zero power

OpenStudy (anonymous):

\[7^2=49\]the other answers are 0, 16, and 1, respectively.

OpenStudy (anonymous):

0^(-4)=0 2^4=16 7^2=49 10^0=1

OpenStudy (anonymous):

So which value is the greatest ?